### v2

# Follow-Bench: A Unified Motion Planning Benchmark for Socially-Aware Robot Person Following

Abstract
Robot person following (RPF)—mobile robots that follow and assist a specific person—has emerging applications in personal assistance, security patrols, eldercare, and logistics. To be effective, such robots must follow the target while ensuring safety and comfort for both the target and surrounding people. In this work, we present the first end-to-end study of RPF, which (i) surveys representative scenarios, motion-planning methods, and evaluation metrics with a focus on safety and comfort; (ii) introduces Follow-Bench, a unified benchmark simulating diverse scenarios, including various target trajectory patterns, dynamic-crowd flows, and environmental layouts; and (iii) re-implements six popular RPF planners, ensuring that both safety and comfort are systematically considered. Moreover, we evaluate the two highest-performing planners from our benchmark on a differential-drive robot to provide insights into real-world deployment. Extensive simulation and real-world experiments provide quantitative insights into the safety-comfort trade-offs of existing planners, while revealing open challenges and future research directions. All code, data, and deployment scripts are publicly available at https://follow-bench.github.io/.
I INTRODUCTION
Robot person following (RPF)[1, 2, 3, 4] is a fundamental capability in human-robot interaction, with applications ranging from personal assistance and security to service robotics. RPF is particularly important for tasks such as eldercare and guided tours in complex, dynamic environments, where robots must be capable of localizing[5], identifying[6], and accompanying[7] a target individual in a socially aware manner. To this end, accompanying behaviors should adhere to principles of safety, comfort, politeness, and context-aware responsiveness [8].
Previous surveys on RPF[1, 3, 4] have provided broad overviews of its core components—such as perception, tracking, planning, control, and human-robot interaction—as well as factors[3] that influence human spatial preferences and expectations. Building on these foundations, this paper narrows the scope to the motion-planning dimension of RPF, with a particular emphasis on two critical and widely prioritized criteria: safety and comfort. Here, safety denotes continuous target tracking and reliable collision avoidance, while comfort primarily refers to maintaining appropriate interpersonal distance with respect to both the target and surrounding pedestrians. Guided by these requirements, we comprehensively review representative scenarios, planning methods, and evaluation metrics; propose a general RPF-driven motion-planning pipeline; and re-implement six popular planners to systematically analyze their performance.
To enable systematic benchmarking, we introduce Follow-Bench, a unified benchmark for evaluating RPF planners under diverse conditions, including various target trajectory patterns, pedestrian-flow patterns, and environmental layouts. Unlike existing RPF benchmarks—which focus primarily on perception (e.g., identifying a specific target as in TPT-Bench[9] and Chen’s dataset[10]) or follow high-level language instructions in simplified environments (e.g., Uni-Navid[11] built on Habitat 3.0[12])—Follow-Bench targets the motion-planning capabilities required in realistic, socially interactive settings.
Most existing RPF studies oversimplify human behavior by modeling pedestrians as randomly moving agents, neglecting realistic human interactions, and typically restrict evaluation to domestic or lab-scale environments with only a few pedestrians. Consequently, very few RPF approaches have been systematically evaluated in socially interactive scenarios, for evaluating the motion-planning capabilities for the purpose of accompanying the target person continuously and naturally while avoiding collision with others. For example, Repiso et al.[13, 14] evaluate their planners in a simulator that extends the social force model (SFM)[15] to account for interactions. However, their simulated scenarios are limited in scope and, critically, not released as an open-source unified benchmark.
In this paper, we address these limitations by introducing Follow-Bench, which explicitly emphasizes two critical requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and other pedestrians in socially interactive environments. Unlike independent navigation tasks[2, 16, 17, 18], which focus primarily on collision avoidance, these dual requirements underscore a fundamental trade-off between safety and comfort. Through extensive experiments with re-implemented popular planners in both simulation and real-world scenarios, we demonstrate how existing methods struggle to effectively balance this trade-off. In summary, the main contributions of this paper are:
A comprehensive review of existing RPF-related motion planners, scenarios and evaluation metrics with an emphasis on safety and comfort.
A unified benchmark, implemented in a lightweight simulator with realistic social interactions, for evaluating RPF motion-planning capabilities across diverse target trajectory patterns, pedestrian flows, and environmental layouts.
A systematic analysis and comparison of six re-implemented popular RPF planners in both Follow-Bench and real-world settings, with in-depth evaluations, discussions, and future research directions.
The remainder of the paper is organized as follows: Sec. II reviews existing RPF research with a focus on representative scenarios, evaluation metrics, and motion planners. Sec. III describes Follow-Bench, including the simulator, scenarios, experimental setup, and evaluation protocols. Simulation and real-world results are presented in Sec. IV. Sec. V discusses challenges, limitations, and future directions, and Sec. VI concludes the paper.
II Literature Review
The behaviors of autonomous mobile robots in RPF tasks can vary substantially depending on factors such as the application context, robot type, and the assisted person’s preferences [1, 3]. In this paper, we focus on two critical and quantifiable requirements that are frequently prioritized: safety and comfort.
Safety is an objective requirement, ensuring that the robot avoids collisions while maintaining continuous observation of the target person.
Comfort, in contrast, is more subjective and is typically reflected in the robot’s motion patterns and the relative distances between the robot and the target person and the surrounding pedestrians (i.e., proxemics [19]). Higher-level aspects of comfort are beyond the scope of this review, such as following and approaching behaviors [13], group following [20], or explicit human-robot interaction [21].
This review builds upon and complements prior surveys of RPF [1, 2, 3, 4] and related topics in social navigation [8, 22], but narrows the focus to representative scenarios (Sec. II-A) that challenge safety and comfort, and the evaluation metrics (Sec. II-B) used to assess performance with respect to these requirements. Furthermore, the reviewed methods (Sec. II-C) are analyzed with an emphasis on how they are designed to address these two key requirements.
II-A Scenarios
We review representative RPF scenarios that pose challenges to both safety and comfort, encompassing both real-world and simulated environments. The difficulty of each scenario is influenced by multiple factors, including the layout complexity of the static environment, the intricacy of the target’s trajectory, and the degree of interference from dynamic objects (i.e., people). Accordingly, we categorize RPF scenarios along three dimensions: target trajectory pattern, dynamic-crowd flow and environmental scale and layout, as shown in Table I.
II-A1 Target Trajectory Pattern
Some existing studies simplify the target trajectory pattern by modeling it as the obstacle-avoiding shortest path between predefined start and end points on a map. However, in RPF tasks, the target’s walking behavior and trajectory characteristics play a crucial role in determining the difficulty of following, introducing unique challenges for RPF control algorithms.
The complexity of a trajectory mainly arises from the target’s turning behaviors, which we categorize by difficulty into smooth turns and sharp turns. Smooth turns are typically associated with natural obstacle avoidance or casual walking patterns. These include standard L-shaped turns and square paths [7, 23, 24, 25, 26, 27], circular or figure-eight-shaped trajectories [7, 28], and consecutive S-shaped motions [29, 30, 27]. In contrast, sharp turns, which involve abrupt changes in direction, are particularly challenging for RPF systems. A common example is the U-turn [7, 28, 29, 30, 31, 32], though other forms such as obtuse-angle turns and triangular trajectories [7, 33] also pose significant difficulties. Leisiazar et al. [29] explicitly evaluated their RPF algorithm’s performance under varying target turning radii, highlighting the sensitivity of RPF systems to trajectory geometry.
The aforementioned turning behaviors typically involve simultaneous changes in both orientation and position during walking. However, in real-world scenarios, humans may adjust their orientation to a random direction. They would first reorient themselves—i.e., rotate in place—before proceeding in a different direction. This decoupling of orientation and translational movement has been explicitly considered in several RPF-related studies [29]. A notable distinction is made by Hu et al.[25] between step turns and spin turns, with the latter referring to in-place rotations that are particularly common during sharp directional changes. These behaviors introduce additional challenges for RPF systems, especially in target orientation estimation[34, 25] and following position adjustment to avoid obstructing the target’s intended path [7].
| Dimension | Category and References | |
| Target Trajectory Pattern | Sharp Turn | Triangle trajectory [7, 33] |
| U-Turn [7, 28, 29, 30, 31, 32] | ||
| Smooth Turn | Square; L-Turn [7, 23, 24, 25, 26, 27] | |
| Circle; “8” trajectory [7, 28] | ||
| Several Turns; “S” trajectory [29, 30, 27] | ||
| Rotate then Walk [25, 7, 29] | ||
| Environmental Scale & Layout | Scale | Small (10m*10m) [31] |
| Mid (30m*30m) [35, 36] | ||
| Large (100m*100m) [37, 32] | ||
| Narrow Layouts | Hallway; Corridor [24, 31, 27, 38, 39] | |
| Doorway; Passageway [13, 40] | ||
| Cluttered Obstacles | Columns; Obstacle Course [23, 40] | |
| Semantic Layouts | Intersection [37, 39] | |
| Elevator* | ||
| Structured Environments* | ||
| 3D Topographies | Stairs; Ramps [41]; Uneven Terrains [41, 42] | |
| Dynamic-Crowd Flow | Pedestrian Number | One [38] |
| Mid (2-10) [43, 29, 10] | ||
| Large (10-20) [44] | ||
| Dense Crowds* | ||
| Directionality and Pattern | Parallel Traffic [38, 39, 45, 46] | |
| Cross Traffic[39, 47, 45, 10] | ||
| Stuck [39, 44] | ||
| Perpendicular Traffic* | ||
| Join / Leave a Crowd* | ||
| Random Crowds* | ||
| Motion Policy | SFM / ESFM [45, 48, 44] | |
| ORCA* | ||
| CV [47, 33] | ||
II-A2 Environmental Scale & Layout
The scale and layout of an environment introduce distinct challenges to RPF. In small room-scale settings [31] or medium-sized single-floor spaces [35, 36], occlusions are frequent because visibility is limited and obstacles are dense; yet, once tracking is lost, re-finding the target is relatively straightforward thanks to the confined search area. Conversely, large-scale venues such as shopping malls [32] or open plazas exhibit fewer occlusions but make search and re-identification harder because the target can move towards a much broader region.
Beyond scale, environmental layout adds layers of complexity. Narrow structures—hallways or corridors, for example—force the robot to keep a safe yet functional distance from both the target person and nearby pedestrians while navigating tight clearances [24, 31, 27, 38, 39]. When a wide space funnels into a doorway or passage, the follower often must switch its relative position to maintain line-of-sight and safety [13, 40]. Cluttered obstacles similarly demand a constantly shifting balance between obstacle avoidance and target pursuit [23, 40]. In addition, semantic cues embedded in the layout call for context-sensitive behavior: at intersections, the robot must anticipate cross-traffic and blind corners [37, 39]; in elevators, it should shrink its proxemic radius to respect the tighter personal space [8]; and in highly repetitive patterns—waiting-room bench grids, supermarket aisles, or office cubicles—it can exploit spatial regularity to plan more predictable and efficient paths. Finally, 3D topographical variations, such as ramps, stairs, and uneven terrains, introduce additional challenges related to stable human localization[41], locomotion and safe navigation[42]. Collectively, these factors underscore that successful RPF requires strategies tuned not only to environmental scale but also to the nuanced features of the surrounding layouts.
II-A3 Dynamic-Crowd Flow
The complexity of RPF tasks increases in crowded scenes due to occlusion and safety risks caused by non-target pedestrians. Especially due to human-human interactions, pedestrian future motion is diverse and varied. In existing RPF simulated environments, some methods assume simplified crowd motion with constant velocity models [47, 33] without considering interactions, while some methods simulate the interactions via SFM or Extended SFM [45, 48, 44]. ORCA is another interaction simulation method, widely adopted in social navigation [8, 49]—shows promise for dense crowds, but its application to RPF remains unexplored.
As for the pedestrian number settings, current evaluations primarily target sparse scenarios: Some studies focus on scenarios with only one other person[38] or medium-size groups (2-10 people) [43, 29, 10]. A few studies evaluate RPF performance in crowds with 20 people[44]. Although some scenarios involve many pedestrians, the people density of these scenarios is usually not provided. Besides, evaluations involving dense and dynamic crowds, like cafeterias and crossings, are limited.
Furthermore, directionality and patterns of pedestrian flows, including crossing[39, 47, 45, 10], parallel passing[38, 39, 45, 46], and getting stuck in traffic[39, 44], also impact the robot’s following policy. For instance, the robot should leave a space for the target person to escape from getting stuck in crowds [44]. Apart from the above factors, additional target-crowd interactions should be considered, e,g., accompanying the target to cross a perpendicular traffic, and joining or leaving a crowd.
II-B Evaluation Metrics
Safety and comfort are the two of the most frequently emphasized criteria in RPF research [8, 3]. Picture a target individual moving through an environment populated by static obstacles and other pedestrians. An ideal person-following robot should (i) avoid any collision and continuously observe the target, (ii) move smoothly, and (iii) remain in a socially acceptable zone relative to both the target and surrounding pedestrians. The core metrics discussed below are summarized in Table II.
In terms of safety, one common metric is the success rate of the following task, typically defined such that the task fails if more than one collision occurs [44]. Another key safety-related consideration is target visibility, especially important in scenarios with frequent occlusions. To this end, the target visibility ratio (TVR) is often used, measuring the proportion of time the robot successfully maintains tracking of the target [50, 51, 52]. When the target becomes occluded, the robot is expected to search and re-identify the target. In this context, the search success rate (SR) and the success weighted by path length (SPL) are commonly adopted [53]. SPL jointly considers both the success of person search and the efficiency of the path taken to complete the task.
Comfort is often assessed in terms of motion smoothness, typically evaluated through velocity, acceleration, and jerk [54, 55]. It also depends on the spatial relationship between the robot and the target. Following proxemic theory [19], previous works [33, 7] define a preferred region around the target, as illustrated in Fig. 1. Maintaining a position within the personal zone (0.45-1.2 meters) is generally considered desirable, and researchers often compute the duration the robot remains within this zone. In addition to distance preference, angle preference is also important (e.g., maintaining a side-following position). For instance, Repiso et.al [46] proposes a metric that measures the angular deviation from the preferred following angle.
Social awareness further requires maintaining appropriate distances not only from the target but also from surrounding pedestrians. Studies in social navigation [56, 57, 58] suggest maximizing the minimum distance to nearby individuals. For example, Arena [56] introduces a metric that quantifies the time spent within the private zone (less than 0.45 meters) of any pedestrian, with shorter durations indicating better social compliance. Path length is also frequently used as a proxy for energy efficiency [44, 56].
It is important to note that proxemic-based quantitative metrics are not universal standards [8, 3]. Preferred following positions are highly context-dependent, varying with the specific RPF task (e.g., eldercare, photography, delivery), the robot’s physical form factor, and even the social attributes of the involved humans (e.g., gender, age). Moreover, several studies argue that proxemic zones are better represented as egg-shaped [59], asymmetrical [57], or even dynamic [60], rather than fixed circular regions.
| Category | Metric | Unit | Explanation | |||
| Safety |
% | Runs with no collision | |||
| SPL | - |
| Comfort | Path Length | [m] | Robot path length | |||
| Velocity (avg.) | [m/s] | Robot velocity | ||||
| Acceleration (avg.) | [m/s2] | Robot acceleration | ||||
| Movement Jerk | [m/s3] | Derivation of acceleration | ||||
| Time in personal zone | [s] |
| Time in private zone | [s] |
| Angle Deviation | [deg] |
II-C RPF-oriented Motion Planning
| Reference | Year | Global Map | Obst. Avd. | Proxemics | Active Occ. Avd. | Person Search | Adap. F-P | Local planner |
| Ferrer et al.[61] | 2017 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | SFM |
| Goldhoorn et al.[52] | 2017 | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | Traj. Roll Out[62] |
| Montesdeoc et al.[63] | 2020 | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | PID |
| Dang et al.[64] | 2022 | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | DWA |
| Sekiguchi et al.[7] | 2021 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | MPC |
| Leisiazar et al.[29] | 2025 | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | RL |
| Wang et al.[65] | 2024 | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | MPC |
| Ye et al.[53] | 2025 | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | MPC |
RPF-oriented motion planning must simultaneously account for safety and comfort. It generally follows the standard navigation pipeline—comprising global path planning and local motion planning—while incorporating additional human-aware functions or constraints. A representative framework is illustrated in Fig. 2, which includes both global and local planners. Global planners typically rely on a prior map [48, 52, 46, 61, 32, 66] to estimate the target’s long-term destination and to compute a global path. These methods often emphasize safety, particularly in re-finding the target after occlusion, in order to maintain continuous visibility [52, 66]. However, constructing a reliable prior map is impractical in many real-world scenarios, especially in large-scale or dynamic environments. Given the dynamic nature of RPF, most approaches therefore focus on local planners, which are responsible for generating safe and socially comfortable person-following behaviors. For this reason, our review primarily concentrates on local planning methods. A summary of representative approaches and their attributes is provided in Table III. For simplicity, we hereafter refer to RPF-oriented local motion planners as RPF planners.
Current RPF planners typically navigate to a goal based on the target’s most recent position, which has been demonstrated more human-like and socially compliant than replicating the target’s historical trajectory [67]. The most straightforward strategy is to predefine a relative goal with respect to the target and use a PID controller to reach it [27, 5, 6, 68]. Some methods [27, 32] further incorporate proxemic constraints to determine socially appropriate accompanying positions. Such constraints generally include distance and relative position. The distance constraint [27] is usually defined as a comfortable following distance within the social space (Fig. 1), while position preferences are commonly categorized into frontal [29], side-by-side [32], and back following [27]. Among these, back following is the most prevalent due to its unobtrusive nature. Frontal following is less common since it requires accurate prediction of the target’s future movements while avoiding obstacles, whereas side-by-side following is typically adopted in socially interactive scenarios.
Beyond proxemic constraints, collision avoidance is essential. To address this, more sophisticated local planning techniques have been proposed, including dynamic window approach (DWA)[69, 31, 64], SFM[14, 70], model predictive control (MPC)[24, 7, 33, 71, 72], and reinforcement learning (RL)[73, 47, 29, 44]. For example, Dang et al.[64] enhance the DWA by incorporating the desired person-following direction into the velocity sampling process, along with collision constraints and smoothness considerations. SFM methods model the robot’s movement using attractive forces toward the target (or their destination), repulsive forces to avoid obstacles, and social forces to respect interpersonal distance from pedestrians[70].
Among these methods, MPC-based approaches have recently gained prominence due to their ability to optimize robot actions over a future time horizon while considering dynamic constraints. This enables proactive adjustments for maintaining safe and comfortable following behavior. MPC formulations can naturally integrate explicit constraints (e.g., speed limits, safety margins), and advances in solvers such as CasADi[74] make real-time optimization feasible. For instance, Sekiguchi et al.[7] introduce a nonlinear-MPC-based framework that predicts the target’s future trajectory and integrates uncertainty as constraints, enabling proactive following within personal space without obstructing the target. In addition, they follow Hu et al.[75] to adaptively select the person-following position based on symmetry for avoiding obstruction to the target. In contrast, RL-based methods[73, 47, 29, 44] learn implicit policies from data using reward-driven training in simulation environments. These methods avoid manually defining constraints but are often limited by sim-to-real transfer issues and lack validation in large-scale real-world deployments.
In addition to proxemics compliance and collision avoidance, maintaining target visibility is another critical aspect of RPF safety. Existing methods address this through active planning or person search. Active planning anticipates occlusion events and proactively adjusts the robot’s position to maintain line-of-sight. For example, Yao et al.[26] address corner occlusion by predicting the target’s turn and switching from back-following to side-following via a PD controller. While effective in specific cases, this method lacks flexibility to diverse occlusion scenarios. Wang et al.[65] propose a more flexible MPC-based approach that introduces human-human and human-obstacle occlusion constraints. However, while maximizing visibility, it may compromise proxemic comfort, resulting in overly distant following behaviors.
Person search provides an alternative strategy when the target is lost. The most basic solution is to navigate the robot to the target’s last known position[76]. More advanced approaches leverage trajectory prediction techniques—such as linear regression[77], Bayesian regression[78], and SVM-based models[79]—to estimate the target’s future location. Despite improvements, these methods still struggle with occlusion in complex environments (e.g., U-turn corners). To address this, Ye et al.[53] represent the predicted trajectory probabilistically and construct a belief-guided field to incrementally update the likelihood of the target’s presence. Furthermore, they introduce a fluid-following field for dynamically tracking occluders and an observation-based potential field to overtake them based on their motion patterns, thus addressing dynamic occlusion scenarios more robustly.
III Follow-Bench
III-A Overview
To comprehensively evaluate the local motion planning capabilities of various RPF planners within a unified framework, we present Follow-Bench. This benchmark extends a Python-based 2D navigation simulator[80] by incorporating realistic interaction simulations, as well as human visibility checks. Several features are inspired by existing social navigation simulators[17, 81]. Here, we adopt a lightweight simulator instead of high-fidelity simulators for its simplicity and compatibility with CPU-only execution. Moreover, it is worth noting that the interaction mechanisms in many high-fidelity simulators[56, 18] are also fundamentally 2D-based: they first generate socially-aware 2D human trajectories and subsequently augment them with 3D animations for enhanced visualization.
Follow-Bench offers diverse scenarios (Sec. III-B), including various target trajectory patterns, dynamic-crowd flow, and environmental layouts, where the robot is tasked with following a specific individual while avoiding obstacles and ensuring safety and comfort for both the target and surrounding people. It includes an exhaustively re-implemented RPF planners (Sec. III-C), along with evaluation metrics (Sec. III-D) that emphasize safety and comfort. The benchmark is designed to be highly extensible, allowing for the addition of new maps, obstacles, agents and agents’ trajectories and motion attributes.
III-B Scenarios
| Experiment | Variable Type | Values | Fixed Conditions |
| Target Trajectory | - | - | - |
| Dynamic (H) | Human number | {5, 10, 15, 20, 25, 30} | - |
| Layout (H) | Corridor humans | {4, 12, 20, 28} | 5.6 m width |
| Intersection humans | {2, 10, 18, 26} | 4.8 m width | |
| Doorway humans | {2, 6, 10, 14} | 2.8 m width | |
| Cluttered space humans | {10, 20, 30, 40} | 30 obstacles | |
| Layout (O) | Corridor width | {6.8, 6.2, 5.6, 5.0} m | 20 humans |
| Intersection width | {6.0, 5.4, 4.8, 4.2} m | 18 humans | |
| Doorway width | {3.4, 3.1, 2.8, 2.5} m | 10 humans | |
| Cluttered space obstacles | {10, 20, 30, 40} | 30 humans | |
| Dynamic (F) | Following configuration | Back / Side at {1.0, 1.5, 2.0, 2.5} m | 20 humans |
| Layout (F) | Following configuration | Back / Side at {1.0, 1.5, 2.0, 2.5} m | Corridor: 20 humans; 5.6 m width |
| Intersection: 18 humans; 4.8 m width | |||
| Doorway: 10 humans; 2.8 m width | |||
| Cluttered space: 30 humans; 30 obstacles |
In this benchmark, we construct a diverse set of scenarios spanning variations in target trajectories, environmental layouts, and dynamic crowd patterns, drawing upon representative RPF settings summarized in Table I.
Target trajectory patterns include two-triangle, two-square, U-turn, figure-eight (‘8’) patterns, L-turns with varying angles (30∘, 45∘, and 60∘), and back-and-forth walking (Fig. 3).
Dynamic-Crowd flows cover parallel crossing, perpendicular crossing, circular crossing, and randomly distributed pedestrian flows (Fig. 4).
Environmental layouts comprise corridors, doorways, intersections, and cluttered spaces populated with dynamic pedestrians (Fig. 5).
These scenario types are widely adopted in RPF-related studies for evaluating the adaptive intelligence of planners. Building on them, we also assess different preferred following positions, including back-following [53] and side-following [46] (Fig. 3).
In all scenarios, human trajectories are generated via an A* global planner over predefined waypoints, while avoiding obstacles based on ORCA or SFM. The baseline configuration models each human as an omni-directional dynamic cylinder (radius: , max speed: ) and the robot as a differential-drive platform () equipped with a 2D LiDAR ( range), subject to translational velocity and acceleration limits. The framework is inherently modular: agent geometry, sensing profile, and motion dynamics can be readily reconfigured to support alternative embodiments, including omni-directional, differential-drive, and Ackermann steering systems.
The evaluated scenarios are summarized in Table IV. To quantify the complexity of socially aware scenarios, we adopt the characterization framework of Stratton et al. [16], evaluating planner performance under varying human densities and environmental occupancy conditions (i.e., passageway widths or obstacle numbers). For dynamic-crowd flows, we define six number levels—5, 10, 15, 20, 25, and 30. For environmental layout scenarios, number levels are set individually according to spatial constraints: Corridor (4, 12, 20, 28) with a width; Intersection (2, 10, 18, 26) with a width; Doorway (2, 6, 10, 14) with a doorway width; and Cluttered space (10, 20, 30, 40) with 30 random obstacles. Furthermore, we vary occupancy conditions by either narrowing passageway widths or increasing obstacle numbers: Corridor (6.8, 6.2, 5.6, 5.0) with 20 humans, Intersection (6.0, 5.4, 4.8, 4.2) with 18 humans, Doorway (3.4, 3.1, 2.8, 2.5) with 10 humans, and Cluttered space (40, 30, 20, 10 obstacles) with 30 humans, respectively. The maximum pedestrian densities, obstacle numbers, and minimum passage widths are determined empirically by observing conditions under which simulated pedestrians become prone to congestion.
In addition, we also evaluate the impact of different person-following positions (back-following and side-following, and different following distances—1.0, 1.5, 2.0, 2.5 meters) on planner performance. These variations aim to assess planners’ adaptability to different social preferences and proxemic requirements. Specifically, we evaluate back- and side-following with following distance in the above all scenarios. While under the variation of following distance, the human number of dynamic-crowd flows is set to 20, and environmental layouts are set to Corridor (20 humans, width), Intersection (18 humans, width), Doorway (10 humans, width), and Cluttered space (30 humans, 30 obstacles).
The benchmark is continuously expanded with new scenarios, and the scenario-generation scripts are publicly available, enabling researchers to freely adjust parameters such as human count, corridor width, and overall scenario scale.
III-C Baselines
As reviewed in Sec. II-C and summarized in Table III, numerous RPF planners have been proposed. However, most implementations are not publicly available and, critically, few address obstacle avoidance and proxemics in an integrated manner. To this end, we re-implement a representative set of well-studied RPF planners validated in real-world dynamic environments and explicitly designed to satisfy both safety and comfort requirements. The re-implemented RPF planners include a SFM-based planner [61], a MPC-based planner [80], a DWA-based planner [64], and an enhanced DWA-based planner that incorporates target-trajectory prediction (DWA w/ Traj.).
Proxemic constraints are considered standard practice in recent RPF literature [27, 33, 7]. Accordingly, each planner maintains a “comfortable” following distance of 1.5 m within the target’s personal space. To mitigate target obstruction, we adopt adaptive following-position adjustments similar to [7, 75]. In addition, each baseline is equipped with a standard RPF search module [77, 82, 83] that re-finds the target after loss by navigating to the Kalman-filter-predicted position. Given the prominence of MPC-based approaches, we additionally reproduce two MPC variants: (i) MPC w/ trajectory prediction (MPC w/ Traj.) [7], which guides the robot along a predicted target trajectory inferred from historical motion. This planner halts when the predicted trajectory exhibits high uncertainty, thereby prioritizing safety; and (ii) MPC with dynamic search fields (MPC w/ DS.) [53], which employs a dynamic search field to intelligently re-find the target after occlusion—either by overtaking occluders or by temporarily following them until an overtaking opportunity arises.
The current release of Follow-Bench focuses on these map-independent local planners, evaluating their fundamental ability to avoid collisions and stably follow the target across diverse environments. To decouple motion-planning evaluation from perception, we assume perfect knowledge of the target and pedestrian positions whenever visible, with occlusion handled by the simulator’s detection logic. Static obstacles are detected using a clustering-based method applied to 2D LiDAR input, following [80]. For completeness, each baseline is implemented in both differential-drive and omnidirectional versions.
III-D Evaluation Protocol and Metrics
Using the re-implemented RPF planners described above, we evaluate their performance across the benchmark scenarios introduced in Sec. III-B. The simulation timestep is fixed at 0.1 s, meaning each step corresponds to 0.1 seconds of simulated time. For each scenario, we first determine the minimum number of simulation steps, , required for the target person to complete their trajectory without the robot present, while accounting for interactions with other pedestrians. Each scenario is then simulated with the robot included for a total of steps—a duration that ensures the target has ample time to finish the trajectory.
Each scenario is tested with 100 randomized trials, with surrounding pedestrians’ trajectories and any static obstacles varied across trials. The robot is always initialized at the preferred following position relative to the target. For each trial, we compute the evaluation metrics listed in Table II. Specifically, search success is defined as the robot successfully re-finding the target within steps after each target loss until the trial ends. A successful trial is defined as one that completes without any collisions and in which the robot recovers tracking within steps of the target’s completion.
IV Analysis of RPF planners
Based on the proposed benchmark, we conduct a series of experiments to evaluate the performance of re-implemented RPF planners. The experiments are designed to analyze how these planners perform under all scenarios (Sec. IV-A), different person-following distances and positions, human densities and environmental occupancy conditions (Sec. IV-B). Here, we simulate human-object and human-human interaction based on ORCA. For simplicity, we show part of the evaluation metrics here, detailed metric evaluation can be found in the released code. In addition, Appendix includes complete evaluation results of all experiments in each scenario.
IV-A Performance under Different Scenarios
The results are shown in Fig. 6. Overall, MPC-based planners (MPC, MPC w/ Traj., and MPC w/ DS.) achieve superior performance in terms of SR compared to the other methods (SFM, DWA and DWA w/ Traj.). This advantage stems from their receding-horizon optimization under dynamic constraints and the smooth biconvex reformulation of collision constraints [80], which enables more adaptive and anticipatory responses to dynamic environments. Such capability is particularly beneficial in person-following scenarios with continuously changing target positions, moving pedestrians, and cluttered or narrow spaces. This also explains why MPC-based planners outperform others on TinTPerson and TVR: their fast response allows the robot to remain within the target’s social space, thereby ensuring comfortable following behavior.
Among these, the MPC w/ Traj. planner achieves the highest SR, ASR, and TVR across both Dynamic and Layout scenarios, highlighting its effectiveness in maintaining target visibility while avoiding collisions. This advantage likely arises from its use of predicted target trajectories, which enable smoother person-following behavior. Through tracking a trajectory rather than a single point, the planner can generate more flexible collision-avoidance maneuvers when encountering obstacles. Representative examples are illustrated in Fig. 7 (a) and (b). Another notable feature is its braking behavior [7] in situations where the predicted target trajectory carries high uncertainty, which further enhances safety. However, this braking tendency leads to reduced performance in the Trajectories scenario, as certain target paths involve sharp turns (e.g., L-turns and two-triangle patterns) that demand rapid adaptation to a safe following position rather than stopping. This reveals an inherent trade-off: while braking improves safety in uncertain environments, it compromises responsiveness to abrupt target maneuvers.
We also observe that MPC w/ DS. performs comparably to the standalone MPC planner across all scenarios. A noteworthy finding is that their SR and ASR are nearly identical in both the Dynamic and Layout scenarios, indicating that almost all failures arise from collisions, leaving little opportunity for the dynamic search field to take effect. This outcome is expected, as our simulated environments involve complex interactions in which pedestrians’ future trajectories are difficult to predict due to frequent and unpredictable encounters. However, the dynamic search field [53] relies heavily on accurate trajectory prediction of occluders to construct its search field. Consequently, when the predicted field is inaccurate, the resulting search actions may increase the risk of collisions. Meanwhile, the MPC planner [80]—which serves as the backbone of MPC w/ DS.—does not incorporate proactive planning for avoiding dynamic occluders. Thus, once target occlusion occurs, the robot has little room to maneuver for overtaking or following occluders, leaving limited opportunity for the dynamic search field to demonstrate its advantages.
IV-B Performance under Different Environmental Settings
The results under different human densities and environmental conditions are shown in Fig. 8. In Dynamic and Layout scenarios, back-following consistently outperforms side-following across all planners and metrics. Back-following reduces interference with pedestrian flows, thereby lowering collision risks and improving target visibility. In contrast, side-following is more prone to occlusions and conflicts with pedestrians or static obstacles, leading to reduced SR, ASR, and TVR. Illustrative examples are shown in Fig. 7 (c) and (d), where side-following trajectories significantly interact with obstacles and pedestrian movements, ultimately leading to collisions. However, it is worth noting that side-following is often perceived as more comfortable and socially acceptable in real-world applications, as it allows for better interaction and communication with the target person [46, 20, 61].
Planner performance degrades with increasing human density, as frequent pedestrian interactions produce less predictable trajectories and reduce available free space, thereby elevating the risk of collisions and occlusions. A similar trend is observed in Layout (O): narrower passageways or higher obstacle counts effectively reduce navigable space, analogous to the effect of increased human density. However, the degradation in Layout (O) is less pronounced than in Layout (H) under back-following. This can be attributed to structural constraints in environments such as corridors, doorways, and intersections, which limit the degrees of freedom in both target and pedestrian trajectories. As a result, back-following becomes nearly equivalent to tracking the target’s trajectory directly, reducing the likelihood of interactions and occlusions with other pedestrians. In contrast, side-following remains highly sensitive, as the follower is exposed to continuous interactions with nearby pedestrians and obstacles. Consequently, side-following consistently underperforms, with success rates falling below 30% across all occupancy conditions.
The results under different following distances are presented in Fig. 9. Overall, both overly close (1.0 m) and overly distant (2.0 m or 2.5 m) following distances degrade the performance of the top-performing planners (e.g., MPC-based, MPC w/ Traj., and MPC w/ DS.) across different environmental layouts (Layout (F)). In contrast, this degradation is less pronounced in Dynamic (F); in some cases, the planners even perform better (e.g., MPC-based and MPC w/ DS.) because open spaces provide greater maneuvering room for collision avoidance. The additional room also facilitates person-search, though at farther distances the robot tends to lose track of the target more frequently. Conversely, in Layout (F), narrow environments constrain the available search space, resulting in more collisions. Many of these collisions arise from ambiguities in the preferred following distances relative to environmental structures. For example, in corridors or intersections, some preferred following positions may lie close to walls, thereby creating riskier situations.
Another notable observation from Fig. 9 is that MPC w/ Traj. performs worse than MPC-based and MPC w/ DS. when the following distance is 1.0, 2.0, or 2.5 m. We attribute this to its braking behavior: at short distances (e.g., 1.0 m), abrupt braking increases the likelihood of collisions with surrounding pedestrians. At longer distances (e.g., 2.0 m and 2.5 m), tracking a distant trajectory slows down responsiveness, whereas the point-tracking strategies of MPC-based and MPC w/ DS. allow quicker adjustments. Faster response is also advantageous for re-finding the target after occlusion. Between the two, MPC w/ DS. demonstrates superior person-search performance compared to MPC-based, as the enlarged search space benefits the construction of the dynamic search field, thereby enabling safer and more robust recovery behaviors.
| Metric | MPC | MPC w/ Traj. | MPC w/ DS. | SFM | DWA | DWA w/ Traj. |
| Time (ms) | 55.8 47.7 | 107.2 87.3 | 55.8 47.6 | 0.1 0.0 | 4.3 2.7 | 60.3 35.7 |
IV-C Planning Time Comparison
The computational efficiency results (as shown in Table V) indicate a clear trade-off between planning quality and runtime. The SFM and DWA-based planners exhibit extremely low computation times (below 15 ms), ensuring real-time feasibility. However, their limited optimization capability may compromise planning quality. In contrast, the MPC-based planners require substantially more computation (50-100 ms), with MPC w/ Traj. being the most time-consuming. Although this additional cost may hinder high-frequency control, it potentially provides stronger optimization-based guarantees in complex environments.
IV-D Real-world Experiments
The above experiments indicate that the MPC-based and MPC w/ Traj. planners are the most effective for person-following in complex environments. We implemented these two planners on a differential-drive robot (Scout-mini) equipped with a ZED2 camera for people detection and target identification [6], and a MID-360 LiDAR for obstacle detection and localization using FAST-LIO [84]. The onboard computer is an Intel NUC 11 mini PC with a Core i7-1165G7 CPU. The evaluated scenarios (Fig. 10) include: (i) a target walking randomly with abrupt turns in open space; (ii) a target walking in an environment with static obstacles; and (iii) a target walking in an environment with both static obstacles and moving pedestrians. The preferred following distance was set to 1.5 m, with the robot maintaining a back-following position. Visualizations are provided in the attached video.
In the first scenario, both planners were able to follow the target stably and respond quickly to abrupt maneuvers. However, MPC w/ Traj. produced smoother trajectories and exhibited less aggressive braking due to its trajectory-tracking strategy. In the second scenario, both planners generally avoided collisions with static obstacles, but the MPC-based planner occasionally became stuck when the target executed sharp turns behind obstacles. By contrast, MPC w/ Traj. leveraged predicted trajectories to generate wider-arc motions that avoided close encounters with obstacles. Such wide-arc detours, however, can cause temporary target loss under limited camera FoV and may reduce user comfort. Similar patterns were observed in the third scenario with moving pedestrians. In addition, both planners successfully avoided collisions with sudden pedestrian disturbances by reversing or detouring. These maneuvers, though effective, were reactive rather than proactive since pedestrian trajectory prediction was not incorporated into the planners.
V Discussion
We present our findings through experiments conducted on Follow-Bench and in real-world environments, and outline the challenges and future directions of RPF-driven planners (Sec. V-A). Furthermore, we discuss the limitations of our benchmark and outline potential future work (Sec. V-B).
V-A Challenges and Future Directions
V-A1 Balancing Trade-offs between Safety and Comfort
Currently, most RPF planners consider comfort constraints by forcing robots to track a point or a trajectory that is in the social space of the target person. However, this reference point or trajectory is often occluded by surrounding pedestrians or static obstacles, which can lead to potential collisions. This is the contradiction between safety and comfort. Although collision-avoidance modules are typically incorporated, the conflicting objectives of maintaining proximity to the target and avoiding obstacles make it challenging for optimization-based planners to consistently generate feasible trajectories—particularly in dynamic environments where the target moves freely. Our experiments further confirm this difficulty: in cluttered settings such as Layout and Dynamic, collisions and occlusions occur more frequently, thereby reducing the overall success rate of person-following. The degradation is even more pronounced in side-following configurations (a position generally considered more comfortable), where the robot is more exposed to interactions with both pedestrians and obstacles.
Rather than following a fixed point or trajectory anchored to the target’s latest position, a more promising direction is to actively estimate and pursue an adaptive trajectory that accounts for predicted human motions and surrounding static structures. Wang et al. [65] propose such an active-following approach by embedding both visibility maximization and collision avoidance as soft constraints within a nonlinear MPC framework. While effective in principle, this formulation introduces multiple coupled factors into the cost function, leading to highly non-convex optimization problems that are difficult to solve in real time. Moreover, the evaluated environments are relatively simple—containing only a few distractors and structural obstacles—and the codebase is not publicly available for straightforward benchmarking. Potential improvements include convexifying certain constraints (e.g., visibility and collision avoidance) or leveraging sampling-based MPC methods [85] to mitigate computational overhead.
Another viable strategy is to decouple the RPF planning problem into hierarchical subproblems, thereby reducing computational complexity. For instance, RPF-Search [53] addresses person-search in dynamic environments by: (i) sampling and assessing candidate waypoints around occluders, and (ii) selecting the best waypoint for overtaking or following, which is then executed by a local planner. Similarly, the active RPF problem can be structured into a high-level planner that generates candidate accompanied points around the target optimized for visibility and safety, based on predicted pedestrian trajectories and environmental context, and a low-level planner that selects the optimal waypoint and computes a collision-free trajectory towards it.
V-A2 Improving Spatial-temporal Obstacle Representation, Prediction and Planning
Current re-implemented RPF planners represent static obstacles as polygons using DBSCAN clustering [86], enabling efficient planning. This approach works well for discrete obstacles, such as cluttered objects (see Fig. 5). However, we found that it is less accurate for continuous obstacles, such as walls and doorways. Moreover, it may incorrectly merge the target and nearby obstacles into a single large polygon, leading to suboptimal or even erroneous planning. Therefore, exploring a more robust and efficient obstacle representation that can handle environments with both static and dynamic obstacles remains a promising direction. In addition, the safe distance between the robot and obstacles should be adaptive to different human-robot coexistence scenarios—larger distances in cluttered environments and smaller distances in narrow spaces (e.g., corridors and doorways)—to balance target visibility and collision avoidance for successful target following.
As shown in experiments, MPC w/ Traj. outperforms the MPC-only planner, highlighting the advantage of leveraging target trajectory forecasts for smoother following. In this work, we only exploit the target’s future trajectory. A natural extension would be to incorporate the predicted trajectories of surrounding pedestrians, which have demonstrated promising results for collision avoidance in independent social navigation [87]. For this purpose, instead of a simple Kalman filter, more advanced predictors that account for human-human interactions, such as Social-GAN [88], may further improve performance. Moreover, within the RPF setting, it is important to consider not only collision avoidance but also target visibility maximization and proxemic relationships with both the target and surrounding pedestrians, guided by their predicted trajectories.
V-A3 Discovering and Utilizing Environmental Patterns
In addition to proactive planning over a short-term horizon (e.g., several seconds for trajectory prediction and motion planning), leveraging long-term environmental patterns could further benefit RPF in complex environments. For instance, when accompanying a target across a crowded crosswalk (similar to parallel crossing in Fig. 4), rather than aggressively occupying a nearby position, the robot may temporarily follow the pedestrian flow [89] until a safe following position becomes available, resulting in more comfortable and safer behavior. Similarly, when the target moves through a crowd (e.g., circular or random formations in Fig. 4 or intersections in Fig. 5), proactively recognizing the underlying crowd pattern and selecting a less congested path [90] to re-approach the target could improve both safety and comfort, even if it requires a temporary sacrifice of target visibility. Such a high-level planner that captures long-term environmental structures while guiding the low-level RPF planner represents a promising research direction.
V-A4 Towards More Expressive Intermediate Representations
Intermediate representations continue to be a fundamental challenge in RPF research. Our experiments reveal that point-based or trajectory-based references, coupled with obstacle-avoidance modules, often fail to maintain stable following behavior in crowded environments with frequent occlusions and obstructions. In addition to active following and hierarchical approaches (discussed in Sec. V-A1), an alternative solution is to develop more expressive intermediate representations that integrate multiple information sources and constraints to guide the robot’s motion planning.
A common strategy is to use cost maps [91, 92] or potential fields [93, 37], which provide continuous cost fields that encode environmental obstacles. These methods allow the robot to select follow points or plan paths within feasible regions, rather than relying on isolated instantaneous target points or trajectories. However, they primarily encode obstacle presence and not explicit distances, and updating these maps in real time can be computationally expensive. Euclidean Signed Distance Fields (ESDF) offer a complementary approach by providing a continuous, differentiable distance field that explicitly encodes obstacle distances. ESDFs enable the robot to plan smooth and safe trajectories for efficient target tracking while supporting real-time updates in dynamic environments [94, 95]. However, updating ESDFs in highly dynamic environments can still be computationally demanding.
Beyond collision avoidance, effective intermediate representations must also incorporate additional constraints to improve RPF planning, such as target visibility and social compliance. However, fulfilling all these constraints can result in computational burdens and potentially conflicting objectives. Future research could explore more efficient, adaptive, and expressive representations, for example, by leveraging learning-based methods [96] to encode complex constraints and interactions in a compact form.
V-B Limitations and Future Work
V-B1 More Context-rich Environments
Currently, our Follow-Bench encompasses the most representative scenarios in RPF research, including various target trajectory patterns, dynamic-crowd flows, and environmental layouts. Nevertheless, it primarily evaluates basic motion-planning capabilities in terms of safety and comfort. Future extensions could introduce more context-rich environments, such as hospitals, factories, museums, households, and offices. In these settings, the target person may exhibit application-specific behaviors and interactions with the environment (e.g., visiting a patient, collaborating with colleagues, or shopping), while the robot could be assigned complementary task-oriented goals during following (e.g., delivering medicine, carrying luggage, or guiding to a destination). Such scenarios would demand more sophisticated reasoning and decision-making from RPF planners.
In addition, the current benchmark is implemented in a 2D simulator for ease of evaluating and developing RPF planners. A high-fidelity and more realistic 3D simulator (e.g., IsaacSim [97]) would enable a richer evaluation of RPF systems under perceptual challenges, including humans’ trajectory prediction, visual occlusion [53], visual localization from quadruped platforms [41], and semantic understanding [36].
V-B2 More Realistic and Diverse Interaction Simulation
Currently, Follow-Bench incorporates human-human and human-object interactions through the ORCA and SFM models, which are commonly used in social navigation research. Nevertheless, these models fall short in capturing the full complexity of real-world interactions, such as heterogeneous walking speeds, diverse interaction patterns, and varied obstacle-avoidance behaviors. Future extensions could exploit real-world datasets [9, 98] to derive more realistic interaction dynamics, or employ more sophisticated interaction models [99] that better approximate human behavior in dynamic environments. Such enhancements would substantially improve the benchmark’s capacity to evaluate RPF planners under challenging and realistic conditions.
VI Conclusions
Robot person following is a fundamental capability in human-robot interaction, with applications spanning personal assistance, security, service, and industrial robotics. This paper provides a comprehensive and structured survey of existing RPF research, with a particular emphasis on two widely recognized and critical criteria: safety and comfort. Guided by these objectives, we review representative scenarios, evaluation metrics, and motion planners. To enable systematic benchmarking of reviewed planners, we introduce Follow-Bench, a unified benchmark implemented in a lightweight simulator that models socially aware interactions, including diverse target trajectory patterns, dynamic-crowd flows, and environmental layouts. To the best of our knowledge, this is the first benchmark to explicitly highlight two essential requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and pedestrians in socially interactive environments. Follow-Bench allows researchers to develop and evaluate new RPF planners, with flexible support for adding maps, obstacles, agents, metrics, and following strategies. All code and deployment scripts are publicly released.
Using Follow-Bench, we evaluate six popular RPF planners that explicitly incorporate safety and comfort constraints, revealing the difficulty of effectively balancing the two. These findings are further corroborated by real-world experiments on a differential-drive robot. Our results show that (i) MPC-based planners consistently outperform others in dynamic environments, and (ii) an MPC variant that integrates predicted target trajectories (MPC w/ Traj.) achieves the best balance between target visibility and collision avoidance. Nevertheless, all planners degrade significantly in high-density scenarios—particularly in side-following configurations—due to frequent occlusions and complex interactions.
Finally, we discuss open challenges and future research directions for RPF-driven planners, alongside the current limitations of our benchmark. We believe that the comprehensive survey and unified evaluation framework presented in this paper will provide a solid foundation for future research, particularly toward developing socially aware RPF planners that can robustly balance safety and comfort in real-world environments.
APPENDIX
VI-A Baselines
We re-implement six RPF planners that prioritize both safety and comfort. Specifically, to ensure safety, we incorporate a person-search function in all baselines, which navigates to the Kalman-filter-predicted position of the target when it is lost. To mitigate obstruction, we adopt adaptive following-position adjustments similar to those in [7, 75]. For comfort, we set the default following distance to 1.5 meters, a widely accepted comfortable distance, except for the experiments involving different following configurations.
The SFM-based planner [61] uses the navigation point to estimate the attractive force, while repulsive forces from obstacles and other people are used to compute the resultant force for velocity control. The detailed algorithm is presented in Algorithm 1. More details about force calculation can be found in [61]. The MPC-based planner utilizes the RDA-planner [80] to trace the navigation point while avoiding collisions with static obstacles and pedestrians. We chose the RDA-planner for its efficient collision constraint formulation, which takes the form of a smooth bi-convex dual, facilitating subsequent parallelization. The detailed algorithm is shown in Algorithm 2. The MPC w/ Traj. (outlined in Algorithm 3) extends the MPC-based planner by tracing a reference path based on the target’s predicted future trajectory. Additionally, inspired by [7], we account for the uncertainty in the predicted trajectory to avoid aggressive behaviors, such as the robot braking when the uncertainty exceeds a threshold. Specifically, we employ a Kalman filter to predict the target’s future trajectory and measure the uncertainty using the normalized innovation squared value, which quantifies the difference between the actual observation and the filter’s prediction.
The MPC w/ DS. detailed in Algorithm 4, enhances the MPC-based planner by incorporating a dynamic search field [53], enabling safe person-search after occlusions. The field consists of two modes: overtaking and fluid-following. In overtaking mode, repulsive forces penalize proximity to occluders, while attractive forces encourage minimal travel cost and improved visibility of the target. When overtaking costs exceed a predefined threshold, the planner switches to fluid-following mode, where the crowd is modeled as a flow field. Velocity and density fields are constructed from pedestrian trajectory histories, and the robot’s goal is selected by minimizing a cost function that balances robot velocity, local crowd velocity, and density. This encourages the robot to follow regions of lower density and align with pedestrian flow, thereby improving safety and comfort.
The DWA-based planner [64] samples admissible linear and angular velocities within the dynamic window and performs forward simulation for each candidate. Trajectories are evaluated using a cost function that combines goal alignment, obstacle clearance, and velocity preference, and the optimal velocity is selected as detailed in Algorithm 5. However, this method considers only the robot’s immediate next state, failing to fully exploit the predictive capability of forward simulation and often resulting in shortsighted behaviors. The DWA w/ Traj. (outlined in Algorithm 6) enhances the DWA-based planner by tracing a reference path derived from the target’s predicted future trajectory and refining the cost function. In particular, the heading cost integrates short-term and long-term orientation objectives, while the distance cost ensures real-time target tracking. The trajectory prediction module is identical to that used in the MPC w/ Traj. planner.
VI-B Detailed Experimental Results
Here, we detail our experimental results of every single scenario under different human numbers, occupied conditions and following distances. Fig. 11 and Fig. 12 show the performance under varying human numbers in dynamic crowd flows and environmental layouts, respectively. Fig. 13 shows the performance under different environmental occupied conditions in environmental layouts. Fig. 14 and Fig. 15 show the performance under varying following distances in dynamic crowd flows and environmental layouts, respectively. More evaluation results of other evaluation metrics can be found in the released code.